OpenStudy (anonymous):

MEDAL find the zeros of the function. f(x)=x^3-4x^2+x+6

OpenStudy (anonymous):

0=0

OpenStudy (anonymous):

None of the choices are that?

OpenStudy (anonymous):

3 , -1 &2

OpenStudy (anonymous):

x^3-3x^2-x^2+x+6 =x^2(x-3)-(x^2-x-6) =x^2(x-3)-(x^2-3x+2x-6) =x^2(x-3)-(x(x-3)+2(x-3)) =x^2(x-3)-(x-3)(x+2) =(x-3)(x^2-x-2) =(x-3)(x^2-2x+x-2) =(x-3)(x(x-2)+1(x-2)) =(x-3)(x-2)(x+1)

OpenStudy (anonymous):

@baybay88